# California, USA

### v1

@@ Abstract
Sequential recommender systems must model long-range user behavior while operating under strict memory and latency constraints. Transformer-based approaches achieve strong accuracy but suffer from quadratic attention complexity, forcing aggressive truncation of user histories and limiting their practicality for long-horizon modeling. This paper presents HoloMambaRec, a lightweight sequential recommendation architecture that combines holographic reduced representations for attribute-aware embedding with a selective state space encoder for linear-time sequence processing. Item and attribute information are bound using circular convolution, preserving embedding dimensionality while encoding structured metadata. A shallow selective state space backbone, inspired by recent Mamba-style models, enables efficient training and constant-time recurrent inference. Experiments on Amazon Beauty and MovieLens-1M datasets demonstrate that HoloMambaRec consistently outperforms SASRec and achieves competitive performance with GRU4Rec under a constrained 10-epoch training budget, while maintaining substantially lower memory complexity. The design further incorporates forward-compatible mechanisms for temporal bundling and inference-time compression, positioning HoloMambaRec as a practical and extensible alternative for scalable, metadata-aware sequential recommendation.
@@ 1 Introduction
  • We introduce a unified preprocessing and evaluation pipeline and empirically validate the proposed architecture on Amazon Beauty [11] and MovieLens-1M [12], demonstrating consistent gains over SASRec and superior performance to GRU4Rec on MovieLens-1M under a constrained 10-epoch training budget.
@@ 6 Results and Analysis
Table 1 summarizes the final-epoch performance of all models under the fixed 10-epoch training budget. Overall, HoloMambaRec consistently outperforms the Transformer-based SASRec baseline on both datasets and achieves the strongest results on MovieLens-1M, while GRU4Rec remains competitive on the sparser Amazon Beauty dataset.
On Amazon Beauty, GRU4Rec attains the highest HR@10 and NDCG@10, reflecting the strength of recurrent inductive bias in highly sparse and noisy interaction sequences. HoloMambaRec nevertheless improves over SASRec by a clear margin, indicating that selective state space dynamics are more effective than quadratic self-attention under constrained sequence length and training budgets. On MovieLens-1M, HoloMambaRec achieves the best overall performance, improving HR@10 by +0.0338+0.0338 and NDCG@10 by +0.0238+0.0238 relative to SASRec, and outperforming GRU4Rec by an even larger margin.
Table 1: HR@10 and NDCG@10 comparison
@@ 7 Practical Notes
Circular convolution enables attribute-aware item representations while preserving the original embedding dimensionality, ensuring that the encoder’s parameter count and memory footprint remain fixed as additional metadata is introduced. The selective state space scan is executed explicitly over timesteps, which makes the recurrent dynamics transparent and interpretable while retaining linear $\mathcal{O}(L)$ computational complexity. All experiments are conducted using a consistent and reproducible configuration with embedding dimension $d=96$, state size $d_{\mathrm{state}}=16$, and $n_{\mathrm{layers}}{=}2$ for Amazon Beauty and $3$ for MovieLens-1M, optimized with AdamW at a learning rate of $10^{-3}$, batch size $64$, and a fixed training budget of $10$ epochs. Although the architecture supports temporal compression through holographic windowed bundling, this mechanism is disabled in the main experiments to isolate the impact of holographic binding and selective state space dynamics on recommendation quality.

### v2

@@ Abstract
Sequential recommender systems must model long-range user behavior while operating under strict memory and latency constraints. Transformer-based approaches achieve strong accuracy but suffer from quadratic attention complexity, forcing aggressive truncation of user histories and limiting their practicality for long-horizon modeling. This paper presents HoloMambaRec, a lightweight sequential recommendation architecture that combines holographic reduced representations for attribute-aware embedding with a selective state space encoder for linear-time sequence processing. Item and attribute information are bound using circular convolution, preserving embedding dimensionality while encoding structured metadata. A shallow selective state space backbone, inspired by recent Mamba-style models, enables efficient training and constant-time recurrent inference. Experiments on Amazon Beauty and MovieLens-1M under a 10-epoch budget show that HoloMambaRec surpasses SASRec on both datasets, attains state-of-the-art ranking on MovieLens-1M, and trails only GRU4Rec on Amazon Beauty, all while maintaining substantially lower memory complexity. The design further incorporates forward-compatible mechanisms for temporal bundling and inference-time compression, positioning HoloMambaRec as a practical and extensible alternative for scalable, metadata-aware sequential recommendation.
@@ 1 Introduction
  • We introduce a unified preprocessing and evaluation pipeline and empirically validate the proposed architecture on Amazon Beauty [11] and MovieLens-1M [12], demonstrating consistent gains over SASRec and state-of-the-art results on MovieLens-1M under a constrained 10-epoch training budget. A holographic compression variant is analyzed as a negative result and positioned as future work.
@@ 6 Results and Analysis
Table 1 reports the final-epoch performance under the fixed 10-epoch budget. Results use the uncompressed setting (use_compression=False), which reflects the best-performing configuration.
On Amazon Beauty, GRU4Rec remains strongest, but HoloMambaRec surpasses SASRec on both HR and NDCG, indicating that selective state space dynamics handle sparse sequences better than quadratic self-attention. On MovieLens-1M, HoloMambaRec delivers the highest HR@10 and NDCG@10, exceeding SASRec by $24.6\%$ HR and $27.8\%$ NDCG and outperforming GRU4Rec by an even larger margin.
Table 1: HR@10 and NDCG@10 comparison (uncompressed setting)
@@
6.1 Compression analysis: optimization paradox (negative result)
[…]
@@ 7 Practical Notes
Circular convolution enables attribute-aware item representations while preserving the original embedding dimensionality, ensuring that the encoder’s parameter count and memory footprint remain fixed as additional metadata is introduced. The selective state space scan is executed explicitly over timesteps, which makes the recurrent dynamics transparent and interpretable while retaining linear $\mathcal{O}(L)$ computational complexity. All experiments are conducted using a consistent and reproducible configuration with embedding dimension $d=96$, state size $d_{\mathrm{state}}=16$, and $n_{\mathrm{layers}}{=}2$ for Amazon Beauty and $3$ for MovieLens-1M, optimized with AdamW at a learning rate of $10^{-3}$, batch size $64$, and a fixed training budget of $10$ epochs. Temporal compression via holographic bundling is evaluated separately as a negative result: it improves latency and memory but destroys ranking accuracy; thus it remains future work rather than part of the core claim.